OpenStudy (anonymous):

1/x-1 + 1 = 4 how do u solve for the unknown?

OpenStudy (anonymous):

It would be nice to show the steps so i can understand how to do it..thanks

OpenStudy (anonymous):

Assuming that it's 1/(x-1) + 1 = 4 then you subtract 1 from both sides then multiply both sides by x-1 and then add 3 and divide by 3 to get 3/4.

OpenStudy (anonymous):

subtract one and get 1 / (x - 1) = 3 this is the same as (x-1)^-1 = 3 raise the whole thing to the power of -1 and get x - 1 = 1/3 add one x = 4/3

OpenStudy (anonymous):

opps meant to say 4/3